Wei has three types of shares: A, B and C. A shares pay 8%,B shares pay 6% and C shares pay 11% dividends. Wei has twice as much invested in B shares as A shares, and has $50000 invested altogether. The yearly return from the share dividends is $4850. How much is invested in each type of share❔
Answer by mananth(16949)

Wei has three types of shares: A, B and C. A shares pay 8%,B shares pay 6% and C shares pay 11% dividends. Wei has twice as much invested in B shares as A shares, and has $50000 invested altogether. The yearly return from the share dividends is $4850. How much is invested in each type of share❔
A shares ----x
B shares-------2x
C=y
x+2x+y= 50000
3x+y=50000........................(1)
8%(x)+6%(2x)+11%(y) = 4850
multiply by 100
8x+12x+11y=485000
20x+11y=485000.............(1)
solve (1) & (2)
3 x + 1 y = 50,000 .............1
Total value
20 x + 11 y = 485,000 .............2
Eliminate y
multiply (1)by -11
Multiply (2) by 1
-33 x -11 y = -550,000
20 x + 11 y = 485,000
Add the two equations
-13 x = -65,000
/ -13
x = 5,000
plug value of x in (1)
3 x + 1 y = 50,000
15,000 + y = 50,000
y = 50,000 -15,000
y = 35,000
y = 35,000
x= 5000.00 Amount in A shares $
2x= 10000, Amount in B shares $
y= 35000.00 Amount in C Shares $
m.ananth@hotmail.ca